# SymTFT out of equilibrium: from time crystals to braided drives and Floquet codes

### 28 Dec 2023 v1

@@ Abstract
Dual TC order: The above discussion holds beyond $G=\mathbb{Z}_{2}$. However, $G=\mathbb{Z}_{2}$ already offers striking new results. First we note that $W_{i}^{(e)}$ and $W_{\tilde{i}}^{(m)}$ enter the SymTFT on the same footing. This implies bulk “dual TC” features for $U_{\text{F}}^{(e,\langle m\rangle)}$, i.e., the “$0\pi$PM” thus far thought to be only a boundary TC Wilczek (2012); Shapere and Wilczek (2012); Watanabe and Oshikawa (2015); Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020).
We now consider $C_{ij}^{\prime}(t)=\bra{n}W_{\tilde{i}}^{(m)}(t)W_{\tilde{j}}^{(m)}\ket{n}$, where |$\ket{n}=\ket{\{x_{k}\}}$, $X_{k}\ket{n}=x_{k}\ket{n}$ ($x_{k}\in\{-1,1\}$) are PM eigenstates. (For brevity, $X_{k}$ also denotes the LIOMs away from the FP.) Via $W_{\tilde{i}}^{(m)}(t)$, BCs became dynamic. To place $C_{ij}^{\prime}(t)$ in $\mathcal{H}_{\text{dyn}}$, we eliminate link variables via
$W_{\tilde{i}}^{(m)}(t)W_{\tilde{j}}^{(m)}=[U_{\text{F}}^{(e,\langle m\rangle)\dagger}]^{t}W_{ij}^{(m)}[U_{\text{F}}^{(e,\langle m\rangle)\prime}]^{t},$ (3)
where, $U_{\text{F}}^{(e,\langle m\rangle)}\!\!\to U_{\text{F}}^{(e,\langle m\rangle)\prime}$ takes $\smash{\overline{W}}^{(e)}\to-\smash{\overline{W}}^{(e)}$, $h_{j}\to-h_{j}$, $h_{k}^{(\text{int})}\to-h_{k}^{(\text{int})}$ ($k=j-1,j$). For $|i-j|\gg\xi$, we have $C_{ij}^{\prime}(t)=(-1)^{t}C_{ij}^{\prime}$ with $C_{ij}^{\prime}=\bra{n}W_{ij}^{(m)}\ket{n}$ the disorder correlator Fradkin (2017). This is long-ranged in PM eigenstates (signaling ~$\tilde{\mathbb{Z}}_{2}$ breaking Ji and Wen (2020)), thus the period-doubled $C_{ij}^{\prime}(t)$ signals the system’s spatio-temporal ~$\tilde{\mathbb{Z}}_{2}$ TC order. We illustrate $C_{ij}^{\prime}(t)$ numerically in Fig. 2.
The $(-1)^{t}$ in $C_{ij}^{\prime}(t)$ here also links to $\pi$ spectral pairing, now between |$\ket{n}$ and $\smash{W_{\tilde{i}}}^{(m)}\ket{n}$, i.e., between distinct BCs, or, within $\mathcal{H}_{\text{dyn}}$, between $U_{\text{F}}^{(e,\langle m\rangle)}$ and $U_{\text{F}}^{(e,\langle m\rangle)\prime}$ spectra with the same $\smash{\overline{W}}^{(m)}$ eigenvalue. This generalizes the corresponding degeneracy of a static PM.
@@
Drive dualities: The (dual) TC drives with $b\notin\mathcal{A}$ in Eq. (2) are complemented by the static $\mathcal{A}=\langle b\rangle$ drives; for $\mathbb{Z}_{2}$, these together yield $\pi$SG, $0\pi$PM, SG and PM as the four Floquet phases. The SymTFT supplies drive dualities: automorphisms $\sigma,\sigma^{\prime}:\mathcal{A}\to\sigma(\mathcal{A}),b\to\sigma^{\prime}(b)$. Here $\sigma$ implements “static” (Kramers-Wannier-type) dualities, while $\sigma^{\prime}$ are “Floquet dualities” Berdanier et al. (2018), taking e.g., $\leftrightarrow\pi\text{SG}$ SG. Within SymTFT, the former corresponds to a domain wall $D^{(\sigma)}_{L}$ wrapping $B_{\text{dyn}}$ Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022, 2023). The latter is also implemented topologically, via a domain wall tube $D^{(\sigma^{\prime})}_{\circ}$ embracing $\smash{\overline{W}}^{(b)}$ in space-time.
Duality-twisted BCs and braided drives: Next, we discuss duality-twisted BCs Schutz (1993); Grimm (2002); Ho et al. (2015); Aasen et al. (2016, ); Tan et al. ; Mitra et al. (2023); Samanta et al. and how, in driven settings, these may lead to non-Abelian defect braiding. We focus on $G=\mathbb{Z}_{2}$, but the generalization will be clear. Consider a short domain wall $D_{\text{dyn}}^{(\sigma)}$ along $B_{\text{dyn}}$, from $l-1/4$ [with twist defect $T_{1}^{(\sigma)}$] to $l+3/4$ [with $T_{2}^{(\sigma)}$], cf. Fig. 1. We define the BA such that its members ending on $l$ or $\tilde{l}$ cross $D_{\text{dyn}}^{(\sigma)}$. Now we push $T_{1}^{(\sigma)}$ into the 2D bulk, cf. Fig. 3a. This pushes $W_{l-1,l}^{(e)}$ and $W_{l-1,l}^{(m)}$ far from $B_{\text{dyn}}$; we thus view $Z_{l-1}Z_{l}$ and $X_{l}$ as nonlocal operators. Their product, $W_{l-1,l}^{(f)}=Z_{l-1}Y_{l}$, is however local since $\sigma(f)=f$ allows us to pull $W_{l-1,l}^{(f)}$ through $T_{1}^{(\sigma)}$ to $B_{\text{dyn}}$.
@@
To complete duality-twisting $U_{\text{F}}^{(b,\mathcal{A})}$, we view $\smash{\overline{W}}^{(b)}$ as a topological ingredient; thus, although it is not a product of local BA members [by featuring $W_{l-1,l}^{(b)}$], we retain it in the drive. We denote the resulting drive by $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$.
The first striking feature is that $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ and its $f$-twisted pair $U_{\text{F}}^{(b,\mathcal{A}),\sigma,f}$ have exactly $\pi$-paired spectra. This is by $W_{l}^{(f)}\equiv W_{\tilde{l}-1}^{(m)}W_{l}^{(e)\dagger}=\tilde{X}_{\tilde{l}-1}Z_{l}$ commuting with all local BA members (upon gauging $Z_{l-1}Y_{l}\to Z_{l-1}\tilde{Z}_{\tilde{l}-1}Y_{l}$ via $Z_{l-1}Z_{l}\to Z_{l-1}\tilde{Z}_{\tilde{l}-1}Z_{l}$), and thus with $H_{0}^{(\mathcal{A})}$ and $H_{1}$, while anticommuting with $\smash{\overline{W}}^{(b)}$ ($b\in\{e,m\}$).
@@
Next, we discuss how making duality-twisting dynamic leads to drives featuring non-Abelian defect braids. We take $D_{\text{dyn}}^{(\sigma)}$ to be present throughout, and denote by $\tilde{D}_{l}^{(\sigma)}$ the act of pushing $T_{1}^{(\sigma)}$ into the 2D bulk (thus duality-twisting BCs). Braids enter via the interplay of $\tilde{D}_{l}^{(\sigma)}$ with $D_{\text{ref}}^{(\sigma)}$; for concreteness we focus on a single exchange $R$ between $T_{1}^{(\sigma)}$ and one of the $D_{\text{ref}}^{(\sigma)}$ twists, cf. Fig. 3. We denote this braided move by $\tilde{D}_{l}^{(\sigma,R)}=R\tilde{D}_{l}^{(\sigma)}$. To study this in Floquet settings, we consider $C_{l}^{(\sigma,R)}(t)=\bra{n}[\tilde{D}_{l}^{(\sigma,R)}]^{-1}(t)\tilde{D}_{l}^{(\sigma,R)}\ket{n}$, featuring the “braided drive” $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=R^{\dagger}U_{\text{F}}^{(b,\mathcal{A}),\sigma}R=i\tilde{\gamma}_{1}U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ (cf. Fig. 3dyn .
@@
Now $U_{\text{F},R}^{(b,\mathcal{A}),\sigma\dagger}\tilde{\gamma}_{q}U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=(-1)^{q}\tilde{\gamma}_{q}$ ($q=1,2$), thus there is an unpaired $\pi$ and zero mode each, both exact. Such unpaired modes are absent in existing TCs. We have $C_{l}^{(\sigma,R)}(t)=\bra{n}[U_{\text{F}}^{(b,\mathcal{A})\dagger}]^{t}[U_{\text{F},R}^{(b,\mathcal{A}),\sigma}]^{t}\ket{n}$; by $\smash{\overline{W}}^{(b)}\tilde{\gamma}_{1}=-\tilde{\gamma}_{1}\smash{\overline{W}}^{(b)}$ ($b\in\{e,m\}$), this toggles between zero and a nonzero value for any eigenstate |$\ket{n}$ of $\smash{\overline{W}}^{(b)}$. Since $\tilde{\gamma}_{1}$ toggles $\smash{\overline{W}}^{(b)}$ for both of $b\in\{e,m\}$, the braided drive pumps $\mathbb{Z}_{2}$ charge between the system and its BCs.
@@
Figure 3: Braid behind the braided drive $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}$ and the corresponding evolution of $\smash{\overline{W}}^{(b)}$ (blue line) through deformations avoiding the twist defects (crosses). The two sides of $\sim$ are related by a contractible loop encircling $T_{1,\text{ref}}^{(\sigma)}$. The sequence shows that $\smash{\overline{W}}^{(b)}$ evolves into $\tilde{\gamma}_{1}\smash{\overline{W}}^{(b)}$ (up to a phase), where the exact zero mode $\tilde{\gamma}_{1}$ corresponds to the noncontractible blue-red loop crossing both $\tilde{D}_{\text{dyn}}^{(\sigma)}$ and $D_{\text{ref}}^{(\sigma)}$.
@@
From TCs to FCs via twisted temporal BCs: Finally we use SymTFT again to further unify BC twisting concepts. First, note that in 2+1D spacetime, inserting $W_{j}^{(b)}$ yields a time-like $b$-anyon worldline at $j\in B_{\text{dyn}}$ [there is no worldline at $B_{\text{ref}}$ since $b$ or $\sigma(b)$ is in $\mathcal{A}_{\text{ref}}$]. Taking time to be periodic, as befits Floquet systems, this gives a noncontractible $W_{j}^{(b)}$ loop wrapping the temporal circle. By spatial $L$-periodicity, this loop appears in space with period $L$. $b$-twisting temporal BCs is the space-time dual of this: the time periodic insertion of noncontractible loops $\smash{\overline{W}}^{(b)}$. Applied to an untwisted drive $U_{\text{F}}^{(\bm{1},\mathcal{A})}$ this precisely yields $\smash{\overline{W}}^{(b)}U_{\text{F}}^{(\bm{1},\mathcal{A})}=U_{\text{F}}^{(b,\mathcal{A})}$. A drive $U_{\text{F}}^{(b,\mathcal{A})}$ thus amounts to $b$-twisted temporal BCs.
@@
What is the space-time dual of duality-twisted BCs? Duality-twisted spatial BCs feature a time-like worldsheet extending from $D_{\text{dyn}}^{(\sigma)}$. The space-time dual is a space-like worldsheet $D_{\text{2D}}^{(\sigma)}$, repeated time periodically. This holds for any automorphism $\sigma$ of a $G$-TO. Automorphism $\sigma$ applied time periodically to anyon worldlines is precisely the topological content of FCs Hastings and Haah (2021); Vuillot ; Davydova et al. (2023); Kesselring et al. ; Vu et al. ; Sullivan et al. (2023); Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. ; these are thus $G$-TO systems with duality-twisted temporal BCs. This viewpoint also extends to FC boundaries; these are such that logical operators (i.e., with ends condensing at $B_{\text{ref}}$ and $B_{\text{dyn}}$) remain logical operators over time Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. ; this implies $\mathcal{A}_{\text{(ref)}}\to\sigma[\mathcal{A}_{\text{(ref)}}]$, as achieved by bending $D_{\text{2D}}^{(\sigma)}$ forward in time at $B_{\text{ref}}$ and $B_{\text{dyn}}$ and continuing it in the future. For a period-$k$ automorphism, $\sigma^{k}=1$, this gives period $k$-tupling at FC boundaries; for $\mathbb{Z}_{2}$ FCs, by $\sigma^{2}=1$ for $e\leftrightarrow m$, this is the recently noted period doubling Aasen et al. .
@@
Our examples used $G=\mathbb{Z}_{2}$, but we described our construction for any finite Abelian $G$. However, while the construction generalizes, going beyond $G=\mathbb{Z}_{2}$ may yield even richer ways to go beyond known Floquet phases, e.g., due to automorphisms spanning a much richer set. Exploring this is just one of the many exciting avenues SymTFT out of equilibrium offers for future research.

### 28 Feb 2024 v2

@@ Abstract
Dual TC order: The above discussion holds beyond $G=\mathbb{Z}_{2}$. However, $G=\mathbb{Z}_{2}$ already offers striking new results. First we note that $W_{i}^{(e)}$ and $W_{i}^{(m)}$ differ only by electric-magnetic duality (cf. Fig. 1). This implies bulk “dual TC” features for $U_{\text{F}}^{(e,\langle m\rangle)}$, i.e., the “$0\pi$PM” thus far thought to be only a boundary TC Wilczek (2012); Shapere and Wilczek (2012); Watanabe and Oshikawa (2015); Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020).
We study these via $C_{ij}^{\prime}(t)=\bra{n}W_{i}^{(m)}(t)W_{j}^{(m)}\ket{n}$, where |$\ket{n}=\ket{\{x_{k}\}}$, $X_{k}\ket{n}=x_{k}\ket{n}$ ($x_{k}\in\{-1,1\}$) are PM eigenstates. (For brevity, $X_{k}$ also denotes the LIOMs away from the FP.) By $W_{i}^{(m)}(t)$, BCs became dynamic. To place $C_{ij}^{\prime}(t)$ in $\mathcal{H}_{\text{dyn}}$, we eliminate link variables via Wde
$W_{i}^{(m)}(t)W_{j}^{(m)}=[U_{\text{F}}^{(e,\langle m\rangle)\dagger}]^{t}W_{ij}^{(m)}[U_{\text{F}}^{(e,\langle m\rangle)\prime}]^{t},$ (3)
where $U_{\text{F}}^{(e,\langle m\rangle)}\!\!\to U_{\text{F}}^{(e,\langle m\rangle)\prime}$ takes $W_{k,k+1}^{(e)}\!\!\to(-1)^{\delta_{jk}}W_{k,k+1}^{(e)}$, hence $h_{j}\to-h_{j}$, $h_{k}^{(\text{int})}\!\!\!\to-h_{k}^{(\text{int})}$ ($k=j-1,j$), and $\smash{\overline{W}}^{(e)}\!\!\to-\smash{\overline{W}}^{(e)}$. For $|i-j|\gg\xi$, we have $C_{ij}^{\prime}(t)=(-1)^{t}C_{ij}^{\prime}$ with $C_{ij}^{\prime}=\bra{n}W_{ij}^{(m)}\ket{n}$. This is long-ranged in PM eigenstates Fradkin (2017) (signaling ~$\tilde{\mathbb{Z}}_{2}$ breaking Ji and Wen (2020)), thus the period-doubled $C_{ij}^{\prime}(t)$ signals spatio-temporal ~$\tilde{\mathbb{Z}}_{2}$ TC order. We illustrate $C_{ij}^{\prime}(t)$ numerically in Fig. 2.
The $(-1)^{t}$ in $C_{ij}^{\prime}(t)$ here also links to $\pi$ spectral pairing, now between |$\ket{n}$ and $\smash{W_{i}^{(m)}}\ket{n}$, i.e., between distinct BCs, or, within $\mathcal{H}_{\text{dyn}}$, between $\smash{U_{\text{F}}^{(e,\langle m\rangle)}}$ and $\smash{U_{\text{F}}^{(e,\langle m\rangle)\prime}}$ spectra with the same $\smash{\overline{W}}^{(m)}$ eigenvalue. This is the dual TC counterpart of the inter-BC degeneracy of a static PM.
@@
Drive dualities: The (dual) TC drives with $b\notin\mathcal{A}$ in Eq. (2) are complemented by the static $\mathcal{A}=\langle b\rangle$ drives; for $\mathbb{Z}_{2}$, these together yield $\pi$SG, $0\pi$PM, SG and PM as the four Floquet phases. The SymTFT supplies drive dualities: automorphisms $\sigma,\sigma^{\prime}:\mathcal{A}\to\sigma(\mathcal{A}),b\to\sigma^{\prime}(b)$. Here $\sigma$ generates “static” (Kramers-Wannier-type) dualities, while $\sigma^{\prime}$ are “Floquet dualities” Berdanier et al. (2018). Within SymTFT, the former corresponds to a domain wall $D^{(\sigma)}_{L}$ wrapping $B_{\text{dyn}}$ Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023a); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022, 2023), yielding a topological, timelike, worldsheet in spacetime. The latter is also implemented topologically, via a domain wall worldsheet $D^{(\sigma^{\prime})}_{\circ}$ embracing $\smash{\overline{W}}^{(b)}$.
Duality-twisted BCs and braided drives: Next, we discuss duality-twisted BCs Schutz (1993); Grimm (2002); Ho et al. (2015); Aasen et al. (2016, ); Tan et al. ; Mitra et al. (2023); Samanta et al. and how, in driven settings, these may lead to non-Abelian defect braiding. We focus on $G=\mathbb{Z}_{2}$, but the generalization will be clear. Consider a short domain wall $D_{\text{dyn}}^{(\sigma)}$ along $B_{\text{dyn}}$, from $l\!-\!1/4$ [with twist defect $T_{1}^{(\sigma)}$] to $l\!+\!3/4$ [with $T_{2}^{(\sigma)}$], cf. Fig. 1. We define the BA such that its members ending on $l$ or $\langle l,l\!+\!1\rangle$ cross $D_{\text{dyn}}^{(\sigma)}$. Now we push $T_{1}^{(\sigma)}$ into the 2D bulk, cf. Fig. 3a. This pushes $W_{l-1,l}^{(e)}$ and $W_{l-1,l}^{(m)}$ far from $B_{\text{dyn}}$ Wde ; thus rendering $Z_{l-1}Z_{l}$ and $X_{l}$ nonlocal. Their product, $W_{l-1,l}^{(f)}=Z_{l-1}Y_{l}$, is however local since $\sigma(f)=f$ allows us to pull $W_{l-1,l}^{(f)}$ through $T_{1}^{(\sigma)}$ to $B_{\text{dyn}}$.
@@
To complete duality-twisting $U_{\text{F}}^{(b,\mathcal{A})}$, we view $\smash{\overline{W}}^{(b)}$ as a topological ingredient; thus, although it is not a product of local BA members for $b\in\{e,m\}$ [by featuring $W_{l-1,l}^{(b)}$], we retain it in the duality-twisted drive, denoted by $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$.
The first striking feature is that $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ and its $f$-twisted pair $U_{\text{F}}^{(b,\mathcal{A}),\sigma,f}\!\!\!=\!W^{(f)\dagger}_{l}U_{\text{F}}^{(b,\mathcal{A}),\sigma}W^{(f)}_{l}$ have exactly $\pi$-paired spectra. This is by $W_{l}^{(f)}=W_{l-1}^{(m)}W_{l}^{(e)\dagger}=\tilde{X}_{l-1}Z_{l}$ commuting with all local BA members (upon gauging $Z_{l-1}Y_{l}\to Z_{l-1}\tilde{Z}_{l-1}Y_{l}$), and thus with $H_{0}^{(\mathcal{A})}$ and $H_{1}$, while anticommuting with $\smash{\overline{W}}^{(b)}$ ($b\in\{e,m\}$).
@@
Next, we discuss how making duality-twisting dynamic leads to drives featuring non-Abelian defect braids. We take $D_{\text{dyn}}^{(\sigma)}$ to be present throughout, and denote by $\tilde{D}_{l}^{(\sigma)}$ the act of pushing $T_{1}^{(\sigma)}$ into the 2D bulk (thus duality-twisting BCs). Braids enter via the interplay of $\tilde{D}_{l}^{(\sigma)}$ with $D_{\text{ref}}^{(\sigma)}$; for concreteness we focus on a single exchange $R$ between $T_{1}^{(\sigma)}$ and one of the $D_{\text{ref}}^{(\sigma)}$ twists, cf. Fig. 3. We denote this braided move by $\tilde{D}_{l}^{(\sigma,R)}=R\tilde{D}_{l}^{(\sigma)}$.
@@
Introducing $[\tilde{D}_{l}^{(\sigma,R)}](t)=[U_{\text{F}}^{(b,\mathcal{A})\dagger}]^{t}[\tilde{D}_{l}^{(\sigma,R)}][U_{\text{F}}^{(b,\mathcal{A})}]^{t}$ we make the braided twist dynamic, which we study via $C_{l}^{(\sigma,R)}(t)=\bra{n}[\tilde{D}_{l}^{(\sigma,R)}]^{-1}(t)\tilde{D}_{l}^{(\sigma,R)}\ket{n}=\bra{n}[U_{\text{F}}^{(b,\mathcal{A})\dagger}]^{t}[U_{\text{F},R}^{(b,\mathcal{A}),\sigma}]^{t}\ket{n}$ dyn . This features the “braided drive” $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=R^{\dagger}U_{\text{F}}^{(b,\mathcal{A}),\sigma}R=i\tilde{\gamma}_{1}U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ (cf. Fig. 3). Now $U_{\text{F},R}^{(b,\mathcal{A}),\sigma\dagger}\tilde{\gamma}_{q}U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=(-1)^{q}\tilde{\gamma}_{q}$ ($q=1,2$): there is an unpaired $\pi$ and zero mode each, both exact. Such unpaired modes are absent in existing TCs.
$\tilde{\gamma}_{1}$ toggles both $\mathbb{Z}_{2}$ and ~$\tilde{\mathbb{Z}}_{2}$ symmetries, exactly, $\smash{\overline{W}}^{(b)}\tilde{\gamma}_{1}=-\tilde{\gamma}_{1}\smash{\overline{W}}^{(b)}$ ($b\in\{e,m\}$). Thus, $C_{l}^{(\sigma,R)}(t)$ vanishes for odd $t$, and is nonzero otherwise, for any $\smash{\overline{W}}^{(b)}$ eigenstate |$\ket{n}$, and the braided drive pumps $\mathbb{Z}_{2}$ charge, exactly, between the system and its BCs.
@@
Figure 3: Braid behind the braided drive $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}$ and the corresponding evolution of $\smash{\overline{W}}^{(b)}$ (blue line) through deformations avoiding the twist defects (crosses). The two sides of $\sim$ are related by a contractible loop encircling $T_{1,\text{ref}}^{(\sigma)}$. The sequence shows that $\smash{\overline{W}}^{(b)}$ evolves into $\tilde{\gamma}_{1}\smash{\overline{W}}^{(b)}$ (up to a phase), where the exact zero mode $\tilde{\gamma}_{1}$ corresponds to the noncontractible blue-red loop crossing both $D_{\text{dyn}}^{(\sigma)}$ and $D_{\text{ref}}^{(\sigma)}$.
@@
From TCs to FCs via twisted temporal BCs: Finally we use SymTFT again to further unify BC twisting concepts. First note that, in 2+1D spacetime, inserting $W_{j}^{(b)}$ Wde yields a timelike $b$-anyon worldline at $j\in B_{\text{dyn}}$ [there is no worldline at $B_{\text{ref}}$ since $b$ or $\sigma(b)$ is in $\mathcal{A}_{\text{ref}}$]. Taking time to be periodic, as befits Floquet systems, this gives a noncontractible $W_{j}^{(b)}$ loop wrapping the temporal circle. “Unfolding” the cylinder in Fig. 1 makes this loop appear with period $L$ in space. $b$-twisting temporal BCs is the space-time dual of this: the time periodic insertion of noncontractible loops $\smash{\overline{W}}^{(b)}$. Applied to an untwisted drive $U_{\text{F}}^{(\bm{1},\mathcal{A})}$ this precisely yields $\smash{\overline{W}}^{(b)}U_{\text{F}}^{(\bm{1},\mathcal{A})}=U_{\text{F}}^{(b,\mathcal{A})}$. A drive $U_{\text{F}}^{(b,\mathcal{A})}$ thus amounts to $b$-twisted temporal BCs.
@@
What is the space-time dual of duality-twisted BCs? Duality-twisted spatial BCs feature a timelike worldsheet extending from $D_{\text{dyn}}^{(\sigma)}$. The space-time dual is a spacelike worldsheet $D_{\text{2D}}^{(\sigma)}$, repeated time periodically. This holds for any automorphism $\sigma$ of a $G$-TO. Automorphism $\sigma$ applied time periodically to anyon worldlines is precisely the topological content of FCs Hastings and Haah (2021); Vuillot ; Davydova et al. (2023); Kesselring et al. ; Vu et al. ; Sullivan et al. (2023); Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. ; these are thus $G$-TO systems with duality-twisted temporal BCs. This viewpoint also extends to FC boundaries; these are such that logical operators (i.e., with ends condensing at $B_{\text{ref}}$ and $B_{\text{dyn}}$) remain logical operators over time Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. ; this implies $\mathcal{A}_{\text{(ref)}}\to\sigma[\mathcal{A}_{\text{(ref)}}]$, as achieved by bending $D_{\text{2D}}^{(\sigma)}$ forward in time at $B_{\text{ref}}$ and $B_{\text{dyn}}$ and continuing it in the future. For a period-$k$ automorphism, $\sigma^{k}=1$, this gives period $k$-tupling at FC boundaries; for $\mathbb{Z}_{2}$ FCs, by $\sigma^{2}=1$ for $e\leftrightarrow m$, this is the recently noted period doubling Aasen et al. .
@@
Our considerations hold for any finite Abelian $G$ and they revealed striking new phenomena already for $G=\mathbb{Z}_{2}$. These may be realized using many-qubit devices akin to those that demonstrated $\mathbb{Z}_{2}$ TCs Mi et al. (2022). Going beyond $G=\mathbb{Z}_{2}$ may yield even richer ways to explore new Floquet phases, e.g., due to automorphisms spanning a much richer set. Exploring this is just one of the many exciting avenues SymTFT out of equilibrium offers for future research.